# On the growth of actions of free products

### v1

@@ 3.4. Topological full groups
Throughout this section we let $Z$ be a Cantor space, and $\mathcal{G}\leq\operatorname{Homeo}(Z)$ be a subgroup of its group of homeomorphisms. Recall that the topological full group of $\mathcal{G}$ is the group $\mathsf{F}(\mathcal{G})$ of all homeomorphisms $h$ of $Z$ such that for every $z\in Z$, there exist a clopen neighbourhood $U$ of $z$ and $g\in\mathcal{G}$ such that $h|_{U}=g|_{U}$. When $\mathcal{G}$ is a cyclic group generated by a single homeomorphism $\varphi$, we write $\mathsf{F}(\varphi)$ instead of $\mathsf{F}(\mathcal{G})$.
We say that an element $g\in\mathsf{F}(\mathcal{G})$ is a 3-cycle if $g^{3}=1$ and the set $\textrm{supp}(g):=\{z\colon g(z)\neq z\}$ is clopen and can be partitioned into 3 distinct clopen subsets $U_{1},U_{2},U_{3}$ such that $g(U_{i})=U_{i+1}$ (with $i$ mod 3). Following Nekrashevych [Nek19], the subgroup of $\mathsf{F}(\mathcal{G})$ generated by 3-cycles is called the alternating full group of $\mathcal{G}$, and is denote $\mathsf{A}(\mathcal{G})$. It is shown in [Nek19] that if the action of $\mathcal{G}$ on $Z$ is minimal (that is, all its orbits are dense), then the group $\mathsf{A}(\mathcal{G})$ is simple and contained in every non-trivial normal subgroup of $\mathsf{F}(\mathcal{G})$. It is also shown in [Nek19] that if $\mathcal{G}$ is finitely generated and its action on $Z$ is expansive and all its orbits have cardinality at least 5, then $\mathsf{A}(\mathcal{G})$ is finitely generated. Recall that an action of a finitely generated group $\mathcal{G}$ on the Cantor space $Z$ is expansive if and only if it is conjugate to a subshift (a $\mathcal{G}$-invariant subset of $\{1,\cdots d\}^{\mathcal{G}}$ for some $d\geq 2$). In many cases, results of Matui show that $\mathsf{A}(\mathcal{G})$ coincides with the commutator subgroup of $\mathsf{F}(\mathcal{G})$ (see his survey [Mat17]).
@@
For a group $G$ acting on $Z$ by homeomorphisms, and for a finite set $Q=\{z_{1},\dots,z_{r}\}\in Z$ we denote by $G_{Q}$ the setwise stabiliser of $Q$, and, and by $G^{0}_{Q}$ its subgroup consisting of elements fix pointwise some neighbourhood of every $z\in Q$. The following is proven in [MB18].
@@
Retain the assumptions of Theorem 3.9, and let $\alpha(n)=\operatorname{vol}_{\mathcal{G},Z}(n)$. Let $X$ be a faithful $G$-set such that $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\alpha(n)^{2}$, and let $H$ be the stabiliser of a point in $X$. Then there exists $z\in Z$ such that $G^{0}_{z}\leq H\leq G_{z}$.
@@
Recall that for any subgroup $G\leq\operatorname{Homeo}(Z)$, the maps $Z\to\operatorname{Sub}(G)$ defined by $z\to G_{z}$ and $z\to G^{0}_{z}$ are respectively upper and lower semi-continuous. When $G$ is countable, semi-continuity implies that $G_{z}=G^{0}_{z}$ for $z$ in a dense $G_{\delta}$ subset.
@@
Assume that $\mathcal{G}$ is finitely generated and its action on $Z$ is minimal and expansive, and let $\alpha(n)=\operatorname{vol}_{\mathcal{G},Z}(n)$. Then for every non-trivial finitely generated group $H$, the group $\mathsf{A}(\mathcal{G})\ast H$ has a Schreier growth gap $n\alpha(n)$.
@@
We shall show that $G$ satisfies the sparse support condition at scale $(\alpha(n),\alpha(n)^{2})$. As usual we fix once and for all a finite symmetric generating set $S$ for $G$, which is used to define all Schreier graphs and distances below. Let $X$ be a faithful $G$-set with $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\alpha(n)^{2}$. We shall show that it satisfies the conclusion of Definition 3.1 with constants $D=C=2$. To this end, fix $R>2$.

### v2

@@ 3.4. Topological full groups
Throughout this section we let $Z$ be a Cantor space, and $\mathcal{G}\leq\operatorname{Homeo}(Z)$ be a subgroup of its group of homeomorphisms. Recall that the topological full group of $\mathcal{G}$ is the group $\mathsf{F}(\mathcal{G})$ of all homeomorphisms $h$ of $Z$ such that for every $z\in Z$, there exist a clopen neighbourhood $U$ of $z$ and $g\in\mathcal{G}$ such that $h|_{U}=g|_{U}$. When $\mathcal{G}$ is a cyclic group generated by a single homeomorphism $\varphi$, we write $\mathsf{F}(\varphi)$ instead of $\mathsf{F}(\mathcal{G})$. The group $\mathsf{F}(\mathcal{G})$ is countable provided $\mathcal{G}$ is countable.
We say that an element $g\in\mathsf{F}(\mathcal{G})$ is a 3-cycle if $g^{3}=1$ and its support $\textrm{supp}(g):=\{z\colon g(z)\neq z\}$ is clopen and can be partitioned into 3 distinct clopen subsets $U_{1},U_{2},U_{3}$ such that $g(U_{i})=U_{i+1}$ (with $i$ mod 3). Following Nekrashevych [Nek19], the subgroup of $\mathsf{F}(\mathcal{G})$ generated by 3-cycles is called the alternating full group of $\mathcal{G}$, and is denoted $\mathsf{A}(\mathcal{G})$. It is shown in [Nek19] that if the action of $\mathcal{G}$ on $Z$ is minimal (that is, all its orbits are dense), then the group $\mathsf{A}(\mathcal{G})$ is simple and contained in every non-trivial normal subgroup of $\mathsf{F}(\mathcal{G})$. It is also shown in [Nek19] that if $\mathcal{G}$ is finitely generated and its action on $Z$ is expansive and all its orbits have cardinality at least 5, then $\mathsf{A}(\mathcal{G})$ is finitely generated. Recall that an action of a finitely generated group $\mathcal{G}$ on the Cantor space $Z$ is expansive if and only if it is conjugate to a subshift (a $\mathcal{G}$-invariant subset of $\{1,\cdots d\}^{\mathcal{G}}$ for some $d\geq 2$). In many cases, results of Matui show that $\mathsf{A}(\mathcal{G})$ coincides with the commutator subgroup of $\mathsf{F}(\mathcal{G})$ (see his survey [Mat17]).
@@
For a group $G$ acting on $Z$ by homeomorphisms, and for a finite set $Q\subset Z$, we denote by $G_{Q}$ the setwise stabiliser of $Q$ and by $G^{0}_{Q}$ its subgroup consisting of elements fix pointwise some neighbourhood of every $z\in Q$. The following is proven in [MB18].
@@
Retain the assumptions of Theorem 3.12, and assume further that $\mathcal{G}$ is finitely generated and that its action on $Z$ is expansive. Let $\alpha(n)=\operatorname{vol}_{\mathcal{G},Z}(n)$. Let $X$ be a faithful $G$-set such that $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\alpha(n)^{2}$, and let $H$ be the stabiliser of a point in $X$. Then there exists $z\in Z$ such that $G^{0}_{z}\leq H\leq G_{z}$.
@@
Recall that for any subgroup $G\leq\operatorname{Homeo}(Z)$, the maps $Z\to\operatorname{Sub}(G)$ defined by $z\mapsto G_{z}$ and $z\mapsto G^{0}_{z}$ are respectively upper and lower semi-continuous. When $G$ is countable, a Baire argument implies that $G_{z}=G^{0}_{z}$ for $z$ in a dense $G_{\delta}$ subset.
@@
Assume that $\mathcal{G}$ is finitely generated and its action on $Z$ is minimal and expansive, and let $\alpha(n)=\operatorname{vol}_{\mathcal{G},Z}(n)$. Then for every non-trivial finitely generated group $H$, the group $G=\mathsf{A}(\mathcal{G})\ast H$ has a Schreier growth gap $n\alpha(n)$.
@@
We shall show that $G$ satisfies the sparse support condition at scale $(\alpha(n),\alpha(n)^{2})$. As usual, we fix once and for all a finite symmetric generating set $S$ for $G$, which is used to define all Schreier graphs and distances below. Let $X$ be a faithful $G$-set with $\operatorname{vol}_{G,X}(n)\not\succcurlyeq\alpha(n)^{2}$. We shall show that it satisfies the conclusion of Definition 3.1 with constants $D=C=2$. To this end, fix $R>2$.